The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.

PSfrag replacements

T = 0 T near T_{c}

T = 0 T near T_{c} T near T_{c} T near T_{c}

1 0 0 1 1 0 0 1 1 0 0 1 1 0 0 1

k/N

lim_{N→∞} 𝔼[f_{1,k}]

k/N

lim_{N→∞} 𝔼[f_{1,k}]

k/N

lim_{N→∞} 𝔼[f_{1,k}]

k/N

lim_{N→∞} 𝔼[f_{1,k}]

Figure 1. Qualitative behavior of lim_{N→∞} 𝔼[f_{1,k}] as a function of k/N in the noninteracting (upper left) and non-spatial large-α Ewens-interacting (upper right) cases. Lower left and lower right show the behavior for finite N.

The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs. The quick brown fox jumped over the lazy dogs.